How do calculate formal charge in a molecule? In detail please.

Grade:12th pass

1 Answers

Jayant Kishore
8 years ago
In case of of the polyatomic ions , the net charge is possessed by an ion as a whole and not by a partcular atom. However for some purposes each atom in a polyatomic molecule/ion is assigned a formal charge. It is basically the difference betwwen the number of valence elctrons in an isolated atom and the no. of electrons assigned to that atom in a Lewis Structure. 
Formal Charge = Total no of valence electrons in free atom – total no. of non bonding (lone pair) electrons 
                          -1/2 total no of bonding (shared ) electrons.
The most stable structure of a molecule is thee one which has smallest formal charges on the atoms.
For example for CS2 ie S=C=S
F.C. on each S atom = valence e- – lone pair e- -1/2 bonding e- = 6 - 4 -1/2 X4     =0
F.C. on C atom = valence e- – lone pair e- -1/2 bonding e-     = 4-0-1/2 X8 =0